OpenStudy (anonymous):

Given the literal equation xy+x/y=2+z, with y = 4x + 2, what is the value of z if x = 1?

OpenStudy (anonymous):

@waterineyes can u please help me u seem really good at math lol

OpenStudy (anonymous):

See, y = 4x + 2 So when x = 1, What will be the y???

OpenStudy (anonymous):

umm i dont know :(... @waterineyes

OpenStudy (anonymous):

Just put x in the equation : y = 4x + 2 y = 4(1) + 2 Now solve it on your own and tell me what is the value of y comes out to be???

OpenStudy (anonymous):

4 1/6?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

What is 4 + 2 = ?????

OpenStudy (anonymous):

6

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Yes, So y = 6 and x = 1 Now put these values in the equation above: 1*6 + 1/6 = 2 + z or, 37/6 - 2 = z z = 25/6 is the answer...